EVA Report – October 25th

EVA Report For 25 October 2018 – EVA3
Author James Burk
Purpose of EVA:
1. Scout Candor Chasma for VR footage
2. Capture 360 and gimbaled camera footage of the MDRS and typical EVA activities.
3. Test out new suits and get feedback.
4. Test out GPS unit.
Destination: Candor Chasma
UTM NAT27 Coordinates: 4250500, 521500
Participants: Max Boyce (Lead), James Burk, Jim Ehrhart, Robert M. Narrative:
We travelled up the road to look for the crossroads to get to Candor but we were not able to find it the first time (or the second time). Eventually we crossed over to the eastbound road and approached the canyon. Around this time we lost connection with Habcom, and stopped our rovers. 2 of us climbed up a ridge (might have been Phobos Peak) to reestablish contact and scout the area. Once we were done, we realized that one of the two rovers was below 60%, and thus we had to immediately return to base.
From the engineering report: Opportunity experienced non-nominal behavior during EVA 3. Rover was carrying larger team members, Robert (5’7”) and James (6’9”) and their knees apparently hit the gear buttons randomly that resulted in the rover being in High gear for portions of the trip and rapid battery consumption. Approximately 100m from the Hab the rover was at 25% and the “motor overheat” lamp came on and the rover stopped. The crew pushed it back to its proper parking spot and used a spare extension cord from the GreenHab to plug it in.

EVA Report For 25 October 2018 – EVA4
Author James Burk
Purpose of EVA:
1. Scout Lith Canyon area for VR footage.
2. Capture 360 and gimbaled camera footage of the MDRS and typical EVA activities.
3. Test out new suits and get feedback.
4. Test out GPS units. Gather GPS waypoint data.
Destination: Lith Canyon
UTM NAT27 Coordinates: 4256046 519257
Participants: Max Boyce (Lead), Shannon Norrell, Marge Lipton, Robert M. Narrative:
All objectives achieved. The team drove the two rovers (Spirit and Curiosity) up to Lith Canyon. As soon as they reached the crossroads north of Pooh’s Corner, they lost Habcom communication as per EVA3. They were able to reach Lith Canyon and survey the area for VR. Also, they took several GPS coordinate readings of various landmarks which we thought future crews might find helpful. We have also converted these Lat/Long readings into NAT27 CONUS.

Mars Sign
Handheld (HH MM SS): 38.24.334 110.47.407
iPhone (Decimal – DDD): 38.4056 110.7903
UTM NAD27 CONUS: 4250840 518292

Pooh’s Corner
Handheld: 38.24.532 110.47.040
iPhone: 38.4081 110.7846
NAD27: 425118 518806

Crossroads Turnoff
Handheld: 38.25.058 110.46.886
iPhone: 38.4175 110.7812
NAD27: 4252162 519101

Route Sign
Handheld: 38.26.437 110.47.792
iPhone: 38.441 110.797
NAD27: 4254766 517716

Burpee
Handheld: 38.27.118 110.47.462
iPhone: 38.4519 110.7909
NAD27: 4256129 517785

Lith Canyon
Handheld: 38.27.146 110.46.767
iPhone: 38.4525 110.7793
NAD27: 4256046 519257

Conversion tools we used:
Convert DDD to Hours, Minutes, Seconds –
https://gisgeography.com/decimal-degrees-dd-minutes-seconds-dms/ Convert – http://www.synnatschke.de/geo-tools/coordinate-converter.php

Journalist Report – October 25th

Journalist’s Report
Marge Lipton, Crew Journalist
Oct 25, 2018

SOL 4: Lots of many small steps before even a tiny leap

Today’s EVA’s were focused on what is still needed before a Virtual Reality training application of the MDRS can show additional terrain via drones. You wouldn’t just want to bring your equipment without doing thorough reconnaissance.

The morning EVA had Jim, James, Robert and Max headed to Candor Chasma. It was estimated to be a 3 hour trip from end to end. While data was being fed back to the Hab, it was noticed that the GPS coordinates (which were being taken for the drone based photogrammetry of terrain for the VR shoot) were not the same as the lat/long on the map in the hab. There are apparently a few different ways of taking coordinates but the videographers are sticking with the GPS readings they personally took on their survey.

Readings were also taken at regular intervals for the battery levels of the rovers, Spirit and Opportunity. Spirit had reasonable readings at the 1 hour 45 minute mark, but Opportunity, had a level of 40% leading the group to turn back. Luckily they did, because yards away from arriving at the Hab a bright red light turned on. Today’s mistake seems to be that putting Opportunity in high gear uses up more juice. So instead of a planned 3 hour EVA, at close to the 2 hour mark, Opportunity needed to be pushed back to the station. It’s a good thing it was almost home by then. The crew is discussing ways it may or may not be possible to bring an extra battery or generator on the rover for just such an emergency in the future.

Robert, Max, Shannon and Marge were on the afternoon EVA. They took Spirit and Curiosity out to the Burpee Dinosaur site. Once there, Shannon took out his drone and flew it over the area to get an idea of what would work. We also took GPS and iPhone coordinates to be used when the actual VR shooting occurs. The group got into the rovers and headed to Lith Canyon after that, but found too much tumbleweed and greenery for it to be Mars.

As for feedback on the space suits, a few us found that while they were fine when walking or sitting still, riding in the rover over some of the potholes caused the helmet to bump into chins and teeth.

Overall our mission has been a success in that it will greatly facilitate the time when the future VR operation occurs.

Sol Summary – October 25th

Crew 197 Sol Summary Report 25Oct2018

Sol 4

Summary Title:

Author’s name: James Burk

Mission Status: Mission is nominal.

Sol Activity Summary: We completed two more EVAs today. The second EVA was very successful trip to Lith Canyon where we flew a drone and took several GPS coordinate readings.

Look Ahead Plan: Tomorrow morning we end Sim, but we will squeeze in one more EVA to scout VR location Candor Chasma which we didn’t reach today.

Anomalies in work: Opportunity rover overheated during our first EVA today (EVA3).

Weather: Dry and clear.

Crew Physical Status: Everybody is great. A little sad we are ending tomorrow because we are really hitting our stride today.

EVAs: Two EVAs accomplished – EVA3 nearly reached Candor Chasma but had to turn back due to a low powered Rover, and one that successfully reached Lith Canyon with all mission objectives accomplished.

Reports to be file:
Sol Summary – James Burk
EVA Report – James Burk
EVA Request – James Burk
Operations Report – Jim Ehrhart
Daily Photos – James Burk
Picture of the Day – James Burk
Journalist Report – Marge Lipton

Support Requested: None

EVA Report Oct 24th

EVA Report

EVA Report For 22 October 2018 – EVA1

Author James Burk

Purpose of EVA:

1. Capture 360 and gimbaled camera footage of the MDRS and typical EVA

activities.

2. Test out new suits and get feedback.

3. EVA and Rover training of crew.

4. Daily engineering readings.

Destination: Pooh’s Corner

UDM27 Coordinates: 38.41 N, 110.78 W

Participants: Max Boyce (Lead), James Burk, Jim Ehrhart, Shannon Norrell.

Narrative:

All objectives achieved. Our 360 and gimballed footage is amazing.

Pooh’s Corner was the original destination, which was quickly achieved

in the first 15 mins of the EVA. We requested permission from Hab

Comm to requisition 2 rovers for a short drive south to Barainca

Butte. On the way, we were intercepted by the director who declined

permission for us to continue the EVA to Barainca Butte. We turned

around and returned to the Hab safely. We used rovers Opportunity and

Curiosity.

EVA Report

EVA Report For 22 October 2018 – EVA2

Author James Burk

Purpose of EVA:

1. Acclimation of additional 4 crewmembers to EVA procedures.

2. Test out new suits and get feedback.

3. Daily engineering readings.

Destination: Pooh’s Corner

UDM27 Coordinates: 38.41 N, 110.78 W

Participants: Susan Jewell (Lead), Sacha Greer, Marge Lipton, Jim

Ehrhart, Robert M.

Narrative:

All objectives achieved.

We used the Spirit rover.

Sol Summary Report Oct 24th

Crew 197 Sol Summary Report 24Oct2018

Sol 3

Summary Title: Think like a Martian, Act like a Martian

Author’s name: James Burk

Mission Status: Mission is nominal.

Sol Activity Summary: We completed our first two EVAs today. Looking

forward to tomorrow’s final day in Sim.

Look Ahead Plan: Two additional EVAs tomorrow to scout VR locations.

Anomalies in work: None

Weather: Dry and a little foggy.

Crew Physical Status: Everybody is great.

EVA: All eight crew members went on EVA to Pooh’s Corner. The first

group of four also attempted to do some additional rover driving but

did not have authorization to proceed south and returned to the Hab.

Reports to be file:

Sol Summary – James Burk

EVA Report – James Burk

EVA Requests – James Burk

Operations Report – Jim Ehrhart

Daily Photos – James Burk

Picture of the Day – James Burk

Journalist Report – Marge Lipton

Support Requested: None

Journalist Report Oct 24th

Sol 3 – Think like a Martian, Act like a Martian

Marge Lipton, Crew Journalist

10/24/2018

After morning yoga and tai chi loosened us up, the first EVA

commenced. Suits were donned, earpieces inserted, radios turned on and

turned in to the correct channel, and suits were fastened with duct

tape at the wrist to keep any Martian air from touching our skin.

EVA #1 was comprised of James, Max, Shannon, and Jim. Before heading

out to Pooh’s Corner, a walk away from the Hab, they checked

engineering, and took photos. Then they split into two groups, two

people took a detour to the Rock Garden, the other two went directly

to Poohs’ Corner. When they all reached that destination they called

in to Hab Com in to find out how much time was left and decided to

come back to base, get a rover and head west for their remaining time.

MISTAKE! When they eventually came back to base, even though the

commander had agreed to their extra excursion, they learned an

important lesson. Think like a Martian, act like a Martian. What that

means is that if you’re given a mission, and you’ve completed it,

don’t take on unnecessary risk. Astronauts on the space station don’t

walk over to the other side when they’ve finished a task just because

they want to. But how are we going to learn if we don’t make mistakes?

Paraphrasing Neil Armstrong in the movie First Man, “Better to fail on

earth than on Mars.” But because our mission is basically about

mapping what needs to be captured in Virtual Reality, that part was a

success.

The next group, Susan, Robert, Sacha, Jim and Marge went out after

lunch to Pooh’s Corner. The protocol is to suit up, gather in the

airlock until being given permission by Hab Com to egress. Our mission

was a success as we wandered around the beautiful Martian surface on

the lookout for interesting rocks and dinosaur bones.

The spacesuit makers, Max and Robert, debriefed us afterwards to hear

what suggestions we had for making them better. Since we’re a photo

mission, besides the comfort and fogginess of some helmets,

consideration was given to where VR and 360 cameras as well as other

paraphernalia could possibly be placed.

Engineering monitored our life supporting equipment. The marvelous HAL

(Habitat Activity Lexica) monitor went down. It’s comprised of various

modules keeping track of crew activity and is in the process of being

revived.

And later tonight we’re having that discussion on the moral dimensions

of space travel. Stay tuned!

Operations Report Oct 24th

  • Crew 197 Operations Report 24Oct18
    SOL: 3
    Name of person filing report: Jim Ehrhart

Non-nominal systems: solar/batteries
Notes on non-nominal systems: batteries apparently on site, to be
installed Friday
Generator (hours run): 23.5, generator on almost full-time, down 5 min
at approx. 5:30pm to add coolant
Solar— SOC % (Before generator is run at night): N/A
Diesel: 75%
Propane – psi: 45
Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons: Not
checked, ATV not in use

Water (trailer) – gallons: N/A trailer not in use
Water (static) – gallons: estimated 90% full
Trailer to Static Pump used – no or yes: No
Water (loft) – Static to Loft Pump used – yes or no: Yes
Water Meter: morning fill, from 10 to 50, meter 138088.98; evening
fill from 10 to 50, meter 138112.92
Toilet tank emptied: no or yes: Yes

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Oil Added? No
ATV Fuel Used: Gals No
# Hours the ATVs were Used today: 0
Notes on ATVs: N/A

Deimos rover used: No, out of service
Hours: 134
Beginning charge: N/A
Ending charge: N/A
Currently charging: No
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Unk
Beginning charge: Unk
Ending charge: Unk
Currently charging: Unk
Spirit rover used: Yes
Hours: 42.3
Beginning charge: 100
Ending charge: 99
Currently charging: Yes
Opportunity rover used: Yes
Hours: 36.1 hrs
Beginning charge: 100
Ending charge: 100
Currently charging: No
Curiosity rover used: Yes
Hours: 55.3 hrs
Beginning charge: 100
Ending charge: 97
Currently charging: Yes

HabCar used and why, where? No

General notes and comments:

Summary of internet: Internet connectivity and bandwidth fine
Summary of suits and radios: All suits performed nominally. One
charger missing and one wonky – will create two new chargers from
existing spare parts tomorrow with suit engineers. Suits experienced
wear and tear on EVAs, some quilting torn in rover ingress/egress, two
scratches on one helmet
All radios reviewed and settings standardized. 13 units in good shape
and fully charged. 4 units not in good condition and set aside as
reserve.
Summary of Hab operations: All systems nominal. Fan/heater air filter
replaced. Loft water filter replaced.
Summary of GreenHab operations: Nominal, smoke detector batteries
replaced and tested
Summary of ScienceDome operations: N/A, not in use due to battery issue
Summary of RAMM operations: Nominal, very disorderly, needs cleaning
and organization
Summary of health and safety issues: Health and Safety Officer perform
checks – see report.
Questions, concerns and requests to Mission Support:
– Battery meter needed to check random spares found in RAMM.
– Suggest light in loft

Creative Report Oct 22nd

Crew 197 – Special Report on HAL system – Sol 1 – October 22, 2018

HAL was a system created by Judd Reed, a former Director of Engineering of the MDRS project. It consisted of a laptop with a custom software stack which the early MDRS crews used to monitor their bandwidth, plan out EVAs, check the weather, and write their crew reports in a standard format. It was also intended to automate the crew report process.

Over the years, HAL fell into disuse because it was not being actively maintain, and at some point it stopped working and the laptop was lost. Judd Reed provided me the source code of HAL and I attempted to recruit volunteers to revise and maintain it. Several people suggested starting fresh with a Raspberry Pi-based solution using modern technologies. RPi is a very low cost linux-based computer that can run all modern applications and software stacks.

Starting in 2017, I began working with Josh Baldwin, Matt Hopson and (later) Ryan Watson to create a new HAL solution, including both hardware and software. What we have today is a custom-built Raspberry Pi powered touchscreen that will be wall-mounted and also has a standard monitor, keyboard, and mouse plugged into it.

Our crew and future crews can use the HAL system to write their daily reports and also for general productivity such as web browsing and office documents. The wall-mounted display is intended to show the most important information that the crew needs to be successful including their current bandwidth usage, their daily schedule, current time (Mountain and UTC) down to the second, current weather
conditions, and finally notifications from MarsChat which is a new time-delayed chat application we are also testing out as part of Crew 197.

An early concept image of HAL is visible on this Marspedia web page: https://marspedia.org/MDRS:HAL_Development_Log

Future crew reports will have a photo or two of the HAL system in operation.

– James Burk, Executive Officer, MDRS Crew 197

Sol Summary October 22nd

Crew 197 Daily Summary Report 23.05.2018

MDRS Daily Summary Report for Sol 1

Summary Title
Morning on Mars

Mission Status:
Good

Sol Activity Summary:
5 of us (James B., Shannon N., Marge, Sacha, and Jim E.) arrived from Grand Junction late yesterday, and were greeted warmly by Dr. Rupert as well as NorCal chapter members Max and Robert who delivered the refurbished suits and helmets. We woke up this morning to our first full day on Mars! We are not yet in sim and there was some cleanup and repairs needed to the tunnels between the Hab and the ancillary facilities. Our commander Dr. Susan Ip-Jewell also arrived this afternoon. After dinner, we had an extended training, briefing, and social session with Dr. Rupert.

Look Ahead Plan: It has rained extensively recently, including today, and it’s pretty muddy outside. We are not expecting to do an EVA until Sol 3 (Wednesday). Tomorrow we will get a full briefing of the suits, and plan to enter sim after lunch. Our mission is a short 5 day mission and we hope to survey future drone flights for the MarsVR Program using EVAs, as well as to ensure all of our crew members have great experiences here.

Anomalies in work:
– Tunnels not fully covered with tarps.
– Science Dome is unusable due to battery failures over the
offseason. New batteries will be installed on Friday.
– Mud outside due to rain, hard to do EVAs.

Weather: Nice in the morning. Rainy in the afternoon but has now cleared up.

Crew Physical Status: Everybody is well and in good spirits.

EVA: None planned for tomorrow.

Reports to be filed:

Operations Report by Dr. Rupert

Sol Summary – James Burk (XO)

Hal Report – James Burk (XO)

Crew Photos – James Burk (XO)

Photo of the Day – James Burk (XO)

Support Requested:

None